# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Acknowledgements.
$U_{5}=\texttt{Rx}(-\pi/2)\texttt{Rz}(-\phi).$ (S24)
S4.1.3 Error Mitigation
@@
Readout Error Mitigation
@@
We then run the algorithm and collect the set of observed bit strings. Since we are interested in the estimation of low-weight observables, we choose to perform mitigation on the marginal distributions. That is, we select the subsystem of the qubit register which is required to estimate the observable, and calculate the bit string distribution of only those qubits. For small subsystems and our experimental value of $10,000$ shots, we are able to generate a statistically accurate bit string distribution. The iterative Bayesian unfolding technique is applied to the subsystem bit string distribution, resulting in a mitigated distribution. The mitigated distribution is then used to calculate the observable value. Because this technique is applied only to low-weight observables in the maximum cut problem (two-point correlations), it is scalable.
Experimental Data
Figure S10: Expectation value of the two-point correlation function $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ for two qubits $i$ and $j$ relevant for building the correlation matrices of all the graph problems based on the QAOA with $p=1$. Expectation values were estimated through the computation of $10^{4}$ bit strings. There are $44$ entries corresponding to $44$ unique subgraphs according to Tab. S3. (a) Raw experimental data from the quantum computer. (b) Readout-error mitigated experimental data. (c) Randomly-compiled experimental data. (d) Readout-error mitigated and randomly compiled experimental data. The Pearson correlation coefficient $r$ is reported. Error bars indicate one standard deviation.
We consider the $44$ expectation values from noiseless state vector emulations of the quantum circuits and from the experimental runs on the superconducting quantum chip Rigetti AnkaaTM-2. There are four experimental scenarios displayed in Fig. S10: (i) raw data, (ii) readout error-mitigated (REM) data, (iii) randomly-compiled (RC) data, and (iv) randomly-compiled and readout error-mitigated data. While differences between the emulation and quantum computer runs are clearly visible, correlations between the expected and experimental data are maintained, as exemplified by the Pearson correlation coefficient $r$ between experimental and emulated expectation values. We observe that error mitigation techniques enhance the correlation between experimental and theoretical data.
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{p,F}=\textrm{tr}\bigl{(}\hat{\rho}_{p,F}\hat{Z}_{i}\hat{Z}_{j}\bigr{)}=F\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{p,F=0}.$ (S25)
@@
S4.2 Classical State Vector Emulations
@@
We have implemented a state vector emulator tailored for the QAOA circuit with $p$ layers of Eq. (S13) using the Python packages NumPy [105] and Numba [106]. Once the normalized quantum state $|\Psi\rangle_{p}$ is obtained, we sample bit strings $\boldsymbol{z}$ according to the probability distribution $|\langle\boldsymbol{z}|\Psi\rangle_{p}|^{2}$.
S5 Additional Data and Analyses
S5.1 Relax-and-Round versus Goemans-Williamson Algorithms
@@
Figure S11: Ratio of the average of various objective values from different solvers as a function of the problem size $N$: $\tilde{C}_{\textrm{GW}}$ is returned by the classical Goemans-Williamson algorithm of Sec. S2.1.8, $\tilde{C}_{\textrm{RR}}$ is returned by the classical relax-and-round algorithm of Sec. S2.1.7, $\tilde{C}_{\textrm{QGW}_{1}}$ is returned by the quantum Goemans-Williamson algorithm with one layer of Sec. S2.2.3, $\tilde{C}_{\textrm{QRR}_{1}}$ is returned by the quantum relax-and-round algorithm with one layer of Sec. S2.2.2. Each data point is averaged over $1,000$ problem instances. Error bars indicate one standard deviation.
@@
We compare the empirical performance of classical and quantum solvers based on a relaxation of the problem: classical relax-and-round (Sec. S2.1.7), classical Goemans-Williamson (Sec. S2.1.8), quantum relax-and-round (Sec. S2.2.2), and quantum Goemans-Williamson (Sec. S2.2.3) algorithms. The main difference between the algorithms is that Goemans-Williamson introduces extra degrees of freedom, such that the objective value of the relaxed version of the problem is tighter to the original binary one. However, whether this translates into a better sign-rounded solution is not guaranteed. We show in Fig. S11 that the two algorithms have similar performances, yielding solutions within $1\%$ of each other. We also observe that for the two algorithms, the classical and quantum counterparts with $p=1$ layer have similar performances. This is explained and discussed further in Sec. S5.5.
@@
For these reasons, we focus on the classical (Sec. S2.1.7) and quantum relax-and-round (Sec. S2.2.2) algorithms.
S5.2 Linear Versus Geometric Temperature Interpolation in Simulated Annealing
@@
Figure S12: Ratio of the average objective values using geometric ($\tilde{C}_{\text{geometric}}$) and linear temperature ($\tilde{C}_{\text{linear}}$) schedules, as a function of the number of simulated annealing sweeps $k$. Each data point is averaged over $1,000$ randomly generated problem instances. Error bars indicate one standard deviation.
@@
In this section we compare the performance of linear and geometric temperature interpolation schemes for interpolating between the hot and cold temperatures used in simulated annealing (See Sec. S2.1.2 and Eqs. (S5) and (S6)). In Fig. S12, we show the ratio of the average objective values using geometric ($\tilde{C}_{\text{geometric}}$) and linear temperature ($\tilde{C}_{\text{linear}}$) schedules, as a function of the number of simulated annealing sweeps $k$. We find that as the problem size $N$ increases ($N\gtrsim 128$), for a number of sweeps ranging from a few tens to a few hundreds, the geometric temperature schedule of Eq. (S6) yields better results. As the number of sweeps $k$ goes to infinity, both temperature schedules should return the same optimal solution, making the ratio converge to one. Throughout this work, we focus on the geometric temperature schedule, unless explicitly specified otherwise.
@@
S5.3 Effect of the Number of Bit Strings for Expectation-Based Quantum Relax-and-Round Algorithm
S5.3.1 Statement
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}\approx\frac{1}{n_{\textrm{ex}}}\sum\nolimits_{k=1}^{n_{\textrm{ex}}}z^{(k)}_{i}z^{(k)}_{j},$ (S26)
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}(n_{\textrm{ex}})\sim\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}(n_{\textrm{ex}}\to+\infty)+\frac{\mathcal{N}(0,1)}{\sqrt{n_{\textrm{ex}}}},$ (S27)
@@
S5.3.2 Numerical Data
Figure S13: (a-d) Average objective value computed using the QRR algorithm based on the QAOA with $p=1$ and $p=2$ layers for different problem sizes $N$, as a function of the number of circuit executions $n_{\textrm{ex}}$ (and, thus, number of bit strings). Correlation matrix elements were computed using Eq. (S26). (a) $N=32$, (b) $N=128$, (c) $N=512$, and (d) $N=2048$. (e-f) Empirical scaling relation for the two-variable average objective value $\tilde{C}(N,n_{\textrm{ex}})$, relating $N$ to $n_{\textrm{ex}}$ following Eq. (S28). (e) $p=1$ with $\kappa=0.2$. (f) $p=2$ with $\kappa=0.3$. Each data point is averaged over $1,000$ problem instances. Error bars indicate one standard deviation.
@@
In the absence of an analytical framework, we perform numerical state vector simulations with $p=1$ and $p=2$ QAOA layers for various problem sizes $N$ and number of circuit executions $n_{\textrm{ex}}$. In Figs. S13a-d, we show the average objective value for different problem sizes $N$, obtained using the QRR algorithm where the correlation matrix elements were estimated using Eq. (S26). We observe that the objective value converges asymptotically with the number of bit strings. We find empirically that the two parameters $N$ and $n_{\textrm{ex}}$ in the average objective value $\tilde{C}(N,n_{\textrm{ex}})$ are connected by a scaling relation
@@
$\tilde{C}\bigl{(}N,n_{\textrm{ex}}\bigr{)}\bigr{/}\tilde{C}\bigl{(}N,n_{\textrm{ex}}\to\infty\bigr{)}=\mathcal{F}\bigl{(}n_{\textrm{ex}}\times N^{-\kappa}\bigr{)},$ (S28)
@@
where $\mathcal{F}$ is a scaling function. Using $\kappa=0.2$ for $p=1$ and $\kappa=0.3$ for $p=2$, respectively, leads to a satisfactory data collapse (Figs. S13e and S13f), supporting the scaling relation of Eq. (S26). Such a scaling relation was also observed in the context of solving Sherrington-Kirkpatrick spin glasses [71] with the QRR algorithm [27]; albeit with a larger exponent $\kappa\approx 1.5$. Thus, at small $p$, the QRR algorithm is much more robust to the number of circuit executions $n_{\textrm{ex}}$ for unit-weight random $3$-regular graphs, which we attribute to the sparser nature of the correlation matrices, compared to Sherrington-Kirkpatrick problem instances. This is consistent with the exponent $\kappa$ increasing with $p$ in Figs. S13e and S13f.
S5.3.3 Implications
Practical meaning
@@
For the QRR algorithm (based on a one-layer QAOA circuit) to maintain a fixed performance on unit-weight random $3$-regular graphs, the number of quantum circuit executions should scale with the problem size as $n_{\textrm{ex}}\sim O(N^{0.2})$. Eq. (S28) and Fig. S13e suggest that $\mathcal{F}(x)\approx 1$ for $x\approx 10^{3}$, i.e., in practice one wants to perform $n_{\textrm{ex}}\gtrsim 10^{3}N^{0.2}$ circuit executions. For the maximum problem size considered $N=4,096$, this means $n_{\textrm{ex}}\gtrsim 5,000$. A similar analysis can be performed at $p=2$ where $n_{\textrm{ex}}\gtrsim 10^{2}N^{0.3}$, yielding $n_{\textrm{ex}}\gtrsim 1,000$ for $N=4,096$. Here, we employ $n_{\textrm{ex}}=10^{4}$ independent of the problem size $N$, ensuring near-optimality of the QRR algorithm.
Algorithmic Complexity of the QRR
@@
Finally, we note that in the absence of a step for finding an efficient swap network on a linear chain topology, encoding $p$ $N$-qubit phase separators takes at most $O(pN)$ layers of parallel two-qubit gates [33]. Because circuit runtime is linear in the number of layers, maintaining a fixed performance for the QRR algorithm requires a runtime of $O(pN^{1+\kappa})$ with $\kappa\ll 1$ for the unit-weight random $3$-regular graphs considered. As such, collecting bit strings might be a limiting factor for the QRR algorithm, at least in terms of algorithmic complexity. In contrast, the eigendecomposition of the correlation matrix has an algorithmic complexity of $O(N)$ (see Sec. S2.2.2). Nevertheless, for larger depth $p$ and intermediate sizes $N$, the eigendecomposition should tend towards an $O(N^{2})$ scaling given that the correlation matrix becomes denser with increasing $p$. The number of bit strings required for asymptotic convergence of the QRR algorithm is unknown in this regime, but one expects that $n_{\textrm{ex}}=1$ bit string be sufficient in the $p\to+\infty$ limit given that it would correspond to the optimal solution. In this situation $\kappa=0$. Therefore, at fixed $N$, we expect a nontrivial relationship between the number of bit strings $n_{\textrm{ex}}$ required to achieve asymptotic convergence, and the depth $p$.
S5.4 Analytical Expression for the Expectation Value of Two-Point Correlations Resulting from the QAOA at $p=1$
Given an $N$-variable binary objective function $\tilde{C}(\boldsymbol{z})=\frac{1}{2}\sum_{i,j=1}^{N}\mathsf{W}_{ij}z_{i}z_{j}$ with $\mathsf{W}_{ij}\in\mathbb{R}$ and $\boldsymbol{z}\in{\pm 1}^{N}$, the quantum state resulting from the QAOA at $p=1$ is [see Sec. S2.2.1]
$\bigl{|}\Psi\bigr{\rangle}=e^{-i\beta\sum_{j=1}^{N}\hat{X}_{j}}e^{i\gamma\hat{\tilde{C}}}\hat{H}^{\otimes N}|{0}\rangle^{\otimes N},$ (S29)

### v2

@@ D.1.2 Compilation
$U_{5}=\texttt{Rx}(-\pi/2)\texttt{Rz}(-\phi).$ (27)
D.1.3 Error Mitigation
@@
Readout Error Mitigation—
@@
We then run the algorithm and collect the set of observed bit strings. Since we are interested in the estimation of low-weight observables, we choose to perform mitigation on the marginal distributions. That is, we select the subsystem of the qubit register which is required to estimate the observable, and calculate the bit string distribution of only those qubits. For small subsystems and our experimental value of $10,000$ shots, we are able to generate a statistically accurate bit string distribution. The iterative Bayesian unfolding technique is applied to the subsystem bit string distribution, resulting in a mitigated distribution. The mitigated distribution is then used to calculate the observable value. Because this technique is applied only to low-weight observables in the maximum-cut problem (two-point correlations), it is scalable.
Experimental Data—
Figure 13: Expectation value of the two-point correlation function $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ for two qubits $i$ and $j$ relevant for building the correlation matrices of all the graph problems based on the QAOA with $p=1$. Expectation values were estimated through the computation of $10^{4}$ bit strings. There are $44$ entries corresponding to $44$ unique subgraphs according to Tab. 3. (a) Raw experimental data from the quantum computer. (b) Readout-error mitigated experimental data. (c) Randomly-compiled experimental data. (d) Readout-error mitigated and randomly compiled experimental data. The Pearson correlation coefficient $r$ is reported. Error bars indicate one standard deviation.
We consider the $44$ expectation values from noiseless state vector emulations of the quantum circuits and from the experimental runs on the superconducting quantum chip Rigetti AnkaaTM-2. There are four experimental scenarios displayed in Fig. 13: (i) raw data, (ii) readout error-mitigated (REM) data, (iii) randomly-compiled (RC) data, and (iv) randomly-compiled and readout error-mitigated data. While differences between the emulation and quantum computer runs are clearly visible, correlations between the expected and experimental data are maintained, as exemplified by the Pearson correlation coefficient $r$ between experimental and emulated expectation values. We observe that error mitigation techniques enhance the correlation between experimental and theoretical data.
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{p,F}=\textrm{tr}\bigl{(}\hat{\rho}_{p,F}\hat{Z}_{i}\hat{Z}_{j}\bigr{)}=F\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{p,F=0}.$ (28)
@@
D.2 Classical State Vector Emulations
@@
We have implemented a state vector emulator tailored for the QAOA circuit with $p$ layers of Eq. (16) using the Python packages NumPy [103] and Numba [104]. Once the normalized quantum state $|\Psi\rangle_{p}$ is obtained, we sample bit strings $\boldsymbol{z}$ according to the probability distribution $|\langle\boldsymbol{z}|\Psi\rangle_{p}|^{2}$.
Appendix E Additional Data and Analyses
E.1 Relax-and-Round versus Goemans-Williamson Algorithms
@@
Figure 14: Ratio of the average of various objective values from different solvers as a function of the problem size $N$: $\tilde{C}_{\textrm{GW}}$ is returned by the classical Goemans-Williamson algorithm of App. B.1.8, $\tilde{C}_{\textrm{RR}}$ is returned by the classical relax-and-round algorithm of App. B.1.7, $\tilde{C}_{\textrm{QGW}_{1}}$ is returned by the quantum Goemans-Williamson algorithm with one layer of App. B.2.3, $\tilde{C}_{\textrm{QRR}_{1}}$ is returned by the quantum relax-and-round algorithm with one layer of App. B.2.2. Each data point is averaged over $1,000$ problem instances. Error bars indicate one standard deviation.
@@
We compare the empirical performance of classical and quantum solvers based on a relaxation of the problem: classical relax-and-round (App. B.1.7), classical Goemans-Williamson (App. B.1.8), quantum relax-and-round (App. B.2.2), and quantum Goemans-Williamson (App. B.2.3) algorithms. The main difference between the algorithms is that Goemans-Williamson introduces extra degrees of freedom, such that the objective value of the relaxed version of the problem is tighter to the original binary one. However, whether this translates into a better sign-rounded solution is not guaranteed. We show in Fig. 14 that the two algorithms have similar performances, yielding solutions within $1\%$ of each other. We also observe that for the two algorithms, the classical and quantum counterparts with $p=1$ layer have similar performances. This is explained and discussed further in App. E.4.
@@
For these reasons, we focus on the classical (App. B.1.7) and quantum relax-and-round (App. B.2.2) algorithms.
E.2 Linear Versus Geometric Temperature Interpolation in Simulated Annealing
@@
Figure 15: Ratio of the average objective values using geometric ($\tilde{C}_{\text{geometric}}$) and linear temperature ($\tilde{C}_{\text{linear}}$) schedules, as a function of the number of simulated annealing sweeps $k$. Each data point is averaged over $1,000$ randomly generated problem instances. Error bars indicate one standard deviation.
@@
In this section we compare the performance of linear and geometric temperature interpolation schemes for interpolating between the hot and cold temperatures used in simulated annealing (See App. B.1.2 and Eqs. (8) and (9)). In Fig. 15, we show the ratio of the average objective values using geometric ($\tilde{C}_{\text{geometric}}$) and linear temperature ($\tilde{C}_{\text{linear}}$) schedules, as a function of the number of simulated annealing sweeps $k$. We find that as the problem size $N$ increases ($N\gtrsim 128$), for a number of sweeps ranging from a few tens to a few hundreds, the geometric temperature schedule of Eq. (9) yields better results. As the number of sweeps $k$ goes to infinity, both temperature schedules should return the same optimal solution, making the ratio converge to one. Throughout this work, we focus on the geometric temperature schedule, unless explicitly specified otherwise.
@@
E.3 Effect of the Number of Bit Strings for Expectation-Based Quantum Relax-and-Round Algorithm
Figure 16: (a-d) Average objective value computed using the QRR algorithm based on the QAOA with $p=1$ and $p=2$ layers for different problem sizes $N$, as a function of the number of circuit executions $n_{\textrm{ex}}$ (and, thus, number of bit strings). Correlation matrix elements were computed using Eq. (29). (a) $N=32$, (b) $N=128$, (c) $N=512$, and (d) $N=2048$. (e-f) Empirical scaling relation for the two-variable average objective value $\tilde{C}(N,n_{\textrm{ex}})$, relating $N$ to $n_{\textrm{ex}}$ following Eq. (31). (e) $p=1$ with $\kappa=0.2$. (f) $p=2$ with $\kappa=0.3$. Each data point is averaged over $1,000$ problem instances. Error bars indicate one standard deviation.
@@ E.3 Effect of the Number of Bit Strings for Expectation-Based Quantum Relax-and-Round Algorithm
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}\approx\frac{1}{n_{\textrm{ex}}}\sum\nolimits_{k=1}^{n_{\textrm{ex}}}z^{(k)}_{i}z^{(k)}_{j},$ (29)
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}(n_{\textrm{ex}})\sim\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}(n_{\textrm{ex}}\to+\infty)+\frac{\mathcal{N}(0,1)}{\sqrt{n_{\textrm{ex}}}},$ (30)
@@
E.3.1 Numerical Data
@@
In the absence of an analytical framework, we perform numerical state vector simulations with $p=1$ and $p=2$ QAOA layers for various problem sizes $N$ and number of circuit executions $n_{\textrm{ex}}$. In Figs. 16a-d, we show the average objective value for different problem sizes $N$, obtained using the QRR algorithm where the correlation matrix elements were estimated using Eq. (29). We observe that the objective value converges asymptotically with the number of bit strings. We find empirically that the two parameters $N$ and $n_{\textrm{ex}}$ in the average objective value $\tilde{C}(N,n_{\textrm{ex}})$ are connected by a scaling relation
@@
$\tilde{C}\bigl{(}N,n_{\textrm{ex}}\bigr{)}\bigr{/}\tilde{C}\bigl{(}N,n_{\textrm{ex}}\to\infty\bigr{)}=\mathcal{F}\bigl{(}n_{\textrm{ex}}\times N^{-\kappa}\bigr{)},$ (31)
@@
where $\mathcal{F}$ is a scaling function. Using $\kappa=0.2$ for $p=1$ and $\kappa=0.3$ for $p=2$, respectively, leads to a satisfactory data collapse (Figs. 16e and 16f), supporting the scaling relation of Eq. (29). Such a scaling relation was also observed in the context of solving Sherrington-Kirkpatrick spin glasses [70] with the QRR algorithm [29]; albeit with a larger exponent $\kappa\approx 1.5$. Thus, at small $p$, the QRR algorithm is much more robust to the number of circuit executions $n_{\textrm{ex}}$ for unit-weight random $3$-regular graphs, which we attribute to the sparser nature of the correlation matrices, compared to Sherrington-Kirkpatrick problem instances. This is consistent with the exponent $\kappa$ increasing with $p$ in Figs. 16e and 16f.
Figure 17: (a) Time in seconds for executing and collecting $n_{\textrm{ex}}$ bit strings from QAOA circuits involving $N$ qubits and $p$ layers, compiled to a linear chain topology using a full swap network [35]. The plot shows the time to execute on the AnkaaTM-2 Quantum Processing Unit accessed through Rigetti Quantum Cloud ServicesTM (QCS), the current circuit duration and the basis for quantum timing in this paper (solid line), and a target execution time based on the best superconducting gate times in the literature. (b) Residuals between the model of Eq. 36 and the experimental data in panel (a).
E.3.2 Implications
Practical meaning—
@@ E.3.2 Implications
For the QRR algorithm (based on a one-layer QAOA circuit) to maintain a fixed performance on unit-weight random $3$-regular graphs, the number of quantum circuit executions should scale with the problem size as $n_{\textrm{ex}}\sim O(N^{0.2})$. Eq. (31) and Fig. 16e suggest that $\mathcal{F}(x)\approx 1$ for $x\approx 10^{3}$, i.e., in practice one wants to perform $n_{\textrm{ex}}\gtrsim 10^{3}N^{0.2}$ circuit executions. For the maximum problem size considered $N=4,096$, this means $n_{\textrm{ex}}\gtrsim 5,000$. A similar analysis can be performed at $p=2$ where $n_{\textrm{ex}}\gtrsim 10^{2}N^{0.3}$, yielding $n_{\textrm{ex}}\gtrsim 1,000$ for $N=4,096$. Here, we employ $n_{\textrm{ex}}=10^{4}$ independent of the problem size $N$, ensuring near-optimality of the QRR algorithm.
Algorithmic Complexity of the QRR—
@@
Finally, we note that in the absence of a step for finding an efficient swap network on a linear chain topology, encoding $p$ $N$-qubit phase separators takes at most $O(pN)$ layers of parallel two-qubit gates [35]. Because circuit run-time is linear in the number of layers, maintaining a fixed performance for the QRR algorithm requires a run-time of $O(pN^{1+\kappa})$ with $\kappa\ll 1$ for the unit-weight random $3$-regular graphs considered. As such, collecting bit strings might be a limiting factor for the QRR algorithm, at least in terms of algorithmic complexity. In contrast, the eigendecomposition of the correlation matrix has an algorithmic complexity of $O(N)$. Nevertheless, for larger depth $p$ and intermediate sizes $N$, the eigendecomposition should tend toward an $O(N^{2})$ scaling given that the correlation matrix becomes denser with increasing $p$. The number of bit strings required for asymptotic convergence of the QRR algorithm is unknown in this regime, but one expects that $n_{\textrm{ex}}=1$ bit string be sufficient in the $p\to+\infty$ limit given that it would correspond to the optimal solution. In this situation $\kappa=0$. Therefore, at fixed $N$, we expect a nontrivial relationship between the number of bit strings $n_{\textrm{ex}}$ required to achieve asymptotic convergence, and the depth $p$.
E.4 Equivalence Between Classical and Quantum Relax-and-Round at $p=1$
E.4.1 Analytical Derivation Based on the QAOA at $p=1$